OpenStudy (rational):

Need help explaining this to my 12 year old niece give one physical example/word problem for each of below : 1) 8/3 2) -8/3 3) 8/-3 4) -8/-3

OpenStudy (rational):

1) 8/3 quotient = 2 remainder = 2 (8) = 2(3) + 2 example : distributing 8 apples to 3 kids each gets 2 apples and 2 is the remainder

OpenStudy (rational):

I am finding it difficult to get models for other 3 problems... please help

OpenStudy (anonymous):

-8 can be the charge of 8 electrons. And they are split among 3 atoms

OpenStudy (rational):

Wow ! nice... that gives me another model to replace charge with debt : distribute a debt of 8 dollars among 3 sons: 2) -8/3 (-8) = -2(3) - 2 quotient = -2, remainder = -2 distribute a debt of 8 dollars among 3 sons each gets a bill of $2, and a debt of $2 still remains

OpenStudy (rational):

two down, two more to go xD

OpenStudy (anonymous):

Since negative integers are no longer natural numbers, things like 'splitting among x' won't work very well.

OpenStudy (rational):

Yeah, I am beginning to see that...

OpenStudy (anonymous):

It's an algebra problem. \(-3x=8\).

OpenStudy (rational):

How many times you need to add -3 to get to 8 ?

OpenStudy (rational):

Yes, converting it to multiplications seems like a good idea !!

OpenStudy (anonymous):

Basically, but obviously you can't count the answer.

OpenStudy (rational):

yeah we can push it to remainder term subtract -3, two times : you will reach 6, the remainder is 2

OpenStudy (rational):

i think i can make it look better... :)

OpenStudy (rational):

need some help on last problem also...

OpenStudy (rational):

3) 8/-3 (8) = -2(-3) + 2 quotient = -2, remainder = 2 example : solve the algebra problem -3x = 8.

OpenStudy (anonymous):

Well, the most common case for this is the slope of a line. The ratio to the vertical change with the horizontal change.

OpenStudy (rational):

beautiful ! xD